2 5 3724042533584 8 | 9 | 5 | 8 | 0 | 2 | 4 | 2 | 6 | 6 | 5|3|6|0|4|||9|||3|||7-8-0-5-4-5-5-9-7-5-9 5 3-1 XI-7-11-797991716861-1181169186918 220 0 7216525773078767376022439 X2815|1313|2|53615351278|6845878718
Answer #1

1. The regression equation is

Y = B0 + B1 X1 + B2 X2 where B0, B1 and B2 are constants and coefficients of Xi's

2. The regression equation is

Y = - 2.3 + 1.88 X1 + 5.15 X2

3.

Given X1= 11 and X2 = 15

The regression equation is

Y = - 2.3 + 1.88 (11) + 5.15 (15) = 95.63

Residual = 75 - 95.63 = -20.63

4.

SSE = 17043

R-square = 0.808

5. 80.8% of variation in the y variable is explained by the independent variable X1 and X2

6. The 95% confdience interval for Y-intercept is (-42.05706898, 37.374361)

The 95% confidence interval for X1 is (1.521200409, 2.233389189)

The 95% confidence interval for x2 is (1.761117505, 8.548867588)